A bar magnet having half magnetic length 6cm calculate the geometric length

Grade:12th

1 Answers

Arun
6 years ago
The magnetic length of a bar magnet is nearly 0.84 times that of the geometric length(Physical length).
 
In question, magnetic length will be = 2*6 = 12 cm.
Hence 
Magnetic length = 0.84 * geometric length
Geometric length = 12/0.84 = 100/7 = 14.28571 cm
 
Regards
Arun (askIITians forum expert)